7.3 Nesmith Corporation's outstanding bonds have a $1,000 par value, a 6% semiannual coupon, 12 years...

7.3

Nesmith Corporation's outstanding bonds have a $1,000 par value, a 6% semiannual coupon, 12 years to maturity, and an 8% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Number of periods = 12 * 2 = 24

Rate = 8% / 2 = 4%

Coupon = (0.06 * 1000) / 2 = 30

Bond price = Coupon *[1- 1 / (1 + r)n] / r + FV/ (1 + r)n

Bond price = 30 * [1- 1 / (1 + 0.04)24] / 0.04 + 1000 / (1 + 0.04)24

Bond price = 30 * [1 - 0.39012] / 0.04 + 390.12147

Bond price = 30 * 15.24696 + 390.12147

Bond price = $847.53